# Open cell property in weakly o-minimal structures

### v1

@@ 3. Weakly o-minimal expansions of ordered groups
Let $\mathcal{M}$ be a non-vaulational weakly o-minimal expansion $\mathcal{M}=(M,<,+,0,\cdots)$ of an ordered group. We say that $\mathcal{M}$ is semisemi-boundedbounded if there exists no definable bijection between a bounded interval and a unbounded interval.
@@
Let $\mathcal{M}$ be a semi-bounded o-minimal expansion $\mathcal{M}=(M,<,+,0,\cdots)$ of an ordered group. Then every non-empty definable open set is a finite union of open cells.
@@
Consider a semi-bounded non-valuational weakly o-minimal expansion $\mathcal{M}=(M,<,+,\dots)$ of an ordered group. Let $U$ be a definable open subset of $M^{n}$. Then there exist finitely many strong open cells whose union is $U$.
@@
Take the smallest refined strong cell decomposition $\mathcal{C}$ of $M^{n}$ partitioning $U$ such that $U=\bigcup_{i=1}^{k}C_{i}$ and $\overline{U}=\bigcup_{i=1}^{k}\overline{C_{i}}$ with $C_{1},\ldots,C_{k}\in\mathcal{C}$. By Remark 2.7, the set $\overline{U}$ is a definable open subset of $\overline{M}^{n}$. Then, by Theorem 3.1, there exist open cells $D_{1},\dots,D_{l}$ such that $\overline{U}=D_{1}\cup\dots\cup D_{l}$. Since each $D_{i}\cap M^{n}$ is an open strong cell and ${U}=(D_{1}\cap M^{n})\cup\dots\cup(D_{l}\cap M^{n})$, this completes the proof. ∎

### v2

@@ 3. Weakly o-minimal expansions of ordered groups
Let $\mathcal{M}=(M,<,+,\cdots)$ be a non-valuational weakly o-minimal expansion of an ordered group. If $\mathcal{M}$} is o-minimal, we say that $\mathcal{M}$ is semi-bounded if there exists no definable bijection between a bounded interval and an unbounded interval [5].
@@
Let $\mathcal{M}=(M,<,+,\cdots)$ be a semi-bounded o-minimal expansion of an ordered group. Then every non-empty definable open set is a finite union of open cells.
@@
Consider a non-valuational weakly o-minimal expansion $\mathcal{M}=(M,<,+,\dots)$ of an ordered group. Let $U$ be a definable open subset of $M^{n}$ and assume that $\overline{{\mathcal{M}}}$ is semi-bounded. Then there exist finitely many open strong cells whose union is $U$.
@@
By Proposition 2.6, there exists an $\overline{{\mathcal{M}}}$-definable open subset $V$ of $\overline{M}^{n}$ such that $U=V\cap M^{n}$. By Theorem 3.1, there exist open cells $D_{1},\dots,D_{l}$ such that $V=D_{1}\cup\dots\cup D_{l}$. Since each $D_{i}\cap M^{n}$ is an open strong cell and ${U}=(D_{1}\cap M^{n})\cup\dots\cup(D_{l}\cap M^{n})$, this completes the proof. ∎
We obtained an improvement of Theorem 3.2 inspired by a private communication with Fujita [8].
Theorem 3.3.
Let $\mathcal{M}=(M,<,+,\dots)$ be a non-valuational weakly o-minimal expansion of an ordered group. Then the following three conditions are equivalent.
(1) For any bounded definable set $I\subseteq M$ and for every $\mathcal{M}$-definable function $f:I\to\overline{M}$, $f(I)$ is bounded.
(2) The canonical o-minimal extension $\overline{\mathcal{M}}$ of $\mathcal{M}$ is semi-bounded.
(3) Any $\mathcal{M}$-definable open set is a finite union of open strong cells.
Proof.
$(1)\Rightarrow(2)$ Assume that (2) does not hold. Then we have a bounded open interval $J_{1}$, an unbounded interval $J_{2}$ and an $\overline{{\mathcal{M}}}$-definable bijection $g:J_{1}\to J_{2}$. Since $\overline{{\mathcal{M}}}$ is o-minimal and by the monotonicity theorem, we may assume $g$ is continuous and strictly monotone. Set $I=J_{1}\cap M$, $f=g|I$. Then $f$ is $\mathcal{M}$-definable. Since $M$ is dense in $\overline{M}$, $f(I)$ is unbounded. This contradicts (1).
$(2)\Rightarrow(1)$ We show the contradiction. Suppose that $I$ is a bounded subset of $M$ and $f:I\to\overline{M}$ is a function definable in $\mathcal{M}$. Assume that $f(I)$ is unbounded. By Proposition 3.1 of [11], we can decompose $I$ into $I=X\cup J_{1}\cup\dots\cup J_{k}$ such that $X$ is finite, each $J_{i}$ is open convex set and $f|J_{i}$ is monotone and strongly continuous. We have the fact that some $f(J_{i})$ is unbounded. Replacing $f$ by $f|J_{i}$, we can assume that $f$ is monotone and strongly continuous. Then the unique extension $\overline{f}:\overline{I}\to\overline{M}$ is monotone and continuous. This is an $\overline{{\mathcal{M}}}$-definable bijection between a bounded open interval $\overline{I}$ and an unbounded open interval $\overline{f}(\overline{I})$. This is a contradiction.
$(2)\Rightarrow(3)$ By Theorem 3.2, this implication follows.
$(3)\Rightarrow(2)$ Assume that (2) does not hold. Then we have an $\overline{{\mathcal{M}}}$-definable bijection $f:J_{1}\to J_{2}$ between a bounded open interval $J_{1}$ and an unbounded interval $J_{2}$. Since $\overline{{\mathcal{M}}}$ is o-minimal and by the monotonicity theorem, we may assume that $f$ is strictly monotone and continuous. We may further assume that $J_{2}$ is bounded below and $f$ is strictly decreasing. By shifting, we can set $J_{1}=(0,b)$. Let $Y$ be the $\overline{{\mathcal{M}}}$-definable open set
$Y=\{(x,y)\in\overline{M}^{2}:(-b<x<b)\land((x<0)\to(y<f(-x)))\land((x>0)\to(y<f(x)))\}$.
Set $X=Y\cap M^{n}$. Then $X$ is $\mathcal{M}$-definable. We prove that $X$ is not a finite union of open strong cells.
Assume $X$ is a finite union of open strong cells. Note that $\{0\}\times M$ is contained in $X$. Thus there exists an open strong cell $C$ and $r\in M$ such that $\{0\}\times(r,\infty)\subseteq C\subseteq X$. Since $\{0\}\times(r,\infty)\subseteq C$, we can find a definable open convex set $I$ with $0\in I$ and a strongly continuous definable function $g:I\to\overline{M}\cup\{-\infty\}$ such that $C=\{(x,y)\in M^{2}:x\in I,y>g(x)\}$. Hence $C$ contains a point $(x_{0},y_{0})$ with $0<x_{0}<b$ and $y_{0}>f(x_{0})$. It contradicts $C\subseteq X$. ∎